Home‎ > ‎Reports‎ > ‎Crew097‎ > ‎

day03


Crew 97 Log Book for January 3, 2011

Photos of the day





Commander's Check-In Report

Judah Epstein Reporting

Crew Physical Status: 
Overall ok. A few crewmembers have minor conditions of a cold. 

Time Departed/Returned from EVA: 
No in-Sim EVAs, but entire crew performed a practice EVA in the preparation for Sim which began tonight. 

Brief Narrative of Field Mission Results: 
Crew practiced suiting up in the EVA suits and going outside of the Hab in the EVA suits. 

EVA Data/Interpretations: 
Crew learned how to perform an EVA. 

Engineering/Hab Maintenance: 
Water to toilet and water to inside potable tank are not refilling with the pumps. The lines are likely frozen due to the extremely cold weather. 

Refill of inside potable tank is conducted via "bucket brigade" and input to toilet is with a bucket of potable water. 

Report Transmission Schedule: 
CMDR Check-in, ENGR, CMDR Report, Photos, Science Report (later in evening) 

Plans for Tomorrow: 
- 07:30 - Breakfast 
- 08:00 - Crew exercise 
- Engineering check-ups - Coordinate EVA radios 
- EVA #1 - Collect pebbles for sedimentary water filter 
- 11:30 - Lunch 
- Engineering - finalize radios, work on ATVs, and water pumps 
- EVA #2 - Collect fine sand for sedimentary water filter 
- Preparation of crew experiments 
- Engineering checks 
- 17:30 - Dinner 
- Reports by 20:00 
- Food Study Questionnaire 
- Showers 

Inventory: 
Food items inventoried by Food Study Coordinator 

Support Requested: 
DG plans bring a new hose for the potable water pump. 

Miscellaneous: 
Need questionnaires for the Food Study (along with specific 2011 Field Season Food Study protocol). 

Crew drove into Hanksville before Sim to acquire final supplies of drinking water, honey, ATV oil from DG, delivered package of bacteria test strips for water filter study, and last ice cream bars until after the Sim!


Commander's Report

Judah Epstein Reporting

Tonight the crew has began the journey of entering the Mars Simulation. 


In preparation for Simulation, the crew conducted final preparations and training. The morning began at 07:30 with breakfast, followed by morning exercise in the Hab. Engineering check-ups were performed before several members of the crew made a stop in Hanksville to acquire final supplies of drinking water, honey, ATV oil from DG, delivered package of bacteria test strips for water filter study, and last ice cream bars until after Sim! 


The crew also trained in EVAs and practiced going outside of the Hab in the EVA suits. 


Experiment preparation was also conducted for final preparations before entering Sim. One crewmember has been working diligently on planning of her physiological study. Commander's water filtration study has begun with greywater influent; and the filter creating effluent of clear water. Water testing is taking place to determine the properties of the filtered water and the amount of success of the sedimentary filtration system. 


Our Engineer has been busy with many engineering tasks and repairs such as water systems, ATVs, determining functionality of the radios, and learning the responsibilities regarding EVA suit maintenance. The water systems are not yet functional, but with the slightly warmer weather perhaps better success will be encountered with less freezing of plumbing lines. One ATV is successfully operating, thus more maintenance will continue on the remaining 3 EVA vehicles. 


Tomorrow will consist of the first full day in Simulation. Two EVAs are planned for tomorrow, consisting of collecting fine sand and pebbles from the MDRS local area for the purpose of creating a third sedimentary filtration system consisting of locally found materials.




Engineering Report

Nathan Wong Reporting

Generator/Electricity: 


Kitty(Diesel Generator) (Used? Yes/No): Yes 


Honey (Gas Generator) (Used? Yes/No): No 


Battery Charge Level: 


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off): 

All systems opperational 



Fuel Status: 


Diesel (Full/0.75/0.5/0.25/Empty): 0.75 


Propane Tank (Full/0.75/0.5/0.25/Empty): 0.66 


Gasoline Tank (Full/0.75/0.5/0.25/Empty): 4 Gas cans full 


Notes/Comments: 



Oil: 


Kitty Oil Quantity (# of Quarts in storage): 3 


Honey Oil Quantity (# of Quarts in storage): 2 


ATV Oil Quantity (# of Quarts in storage): 2 


Car Oil Quantity (# of Quarts in storage): none 



Water Status: 


Outside Potable Water Tank Level (inches from bottom): 40% 


Trailer Potable Water Tank Level (inches from bottom): empty 


Hab Potable Water Tank Level (inches from bottom): 75% 


Potable Water Meter Reading: 40355.7 


In to GreenHab Meter Reading: 


Notes/Comments: 


-Grey Water Flush line is not working. Flush tank is full. Suction is not being generated by the pump, bladder is not damaged. -Pump to Hab tank is not functioning 



GreenHab: 


Condition of plants in Tank 1: N/A 


Condition of plants in Tank 2: N/A 


Condition of plants in Tank 3: N/A 


Greenhab Notes/Comments: Plants not at Hab yet. 



Observatory: 


Telescope Used? (Yes/No): No 


Observatory Notes/Comments: Observatory not used 



Transportation: 


Hab Car (NOT RENTAL) 

Used (yes/no): Yes 

Oil Checked (yes/no): No 

Tire Status: good 

Odometer at end of day: 123085 

Notes/Comments on Hab Car: 


ATV 1 (Opportunity): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 2 (Spirit): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 3 (Viking I): 

Used: No 

Oil Checked: No 

Fuel Consumed: No 

Tire Status: Inflated 

Comments: 


ATV 4 (Viking II): 

Used: Yes 

Oil Checked: No 

Fuel Consumed: Yes 

Tire Status: Inflated 

Comments: 


Viking II was the only operational ATV. Spirit and Opportunity were close to starting, and Viking I was not close to starting. Viking II was run for approximately 10 minutes and refilled with gas. 


The Hab Car was used to get supplies from the Market before starting simulation. 



Heating and Ventilation: 

Thermostat setting upstairs (Farenheit): 67 

Actual temperature upstairs (Farenheit): 75 


Thermostat setting downstairs (Farenheit): 63 

Actual temperature downstairs (Farenheit): 57 



Computers/Networking Infrastructure: 


Notes/comments on internet/computers?: 

Everything is working well 



General Engineering Notes/Comments: 


Everything is working well except the grey water line, and the pump to the Hab tank.



Cook's Report

Nathan Wong Reporting

1) Was today 

- a cooking day? yes 


2) List the foods served at today's main meal (usually the evening meal), giving full names of each. 

Spaghetti with tomato sauce and assorted rehydrated vegetables such as green beans, tomatoes, spinach, peas. 

Can of chili with velveeta cheese tortillas. 

Carrot cake with maple frosting. 


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded if inedible) 

About 4 servings of carrot cake 

Three tortillas 


4) Recipes for anything more complicated than following package directions exactly, or rehydrating. 

Tomato sauce with vegetables – recipe not reported 

Carrot cake – recipe not reported 

maple icing – recipe not reported 


5) Comments/questions 

- any shortages of food supplies or equipment? (includes "cannot find..." ) no 

- any problems with supplied food? (e.g. spoilage, insect or rodent damage, off flavors, difficulty rehydrating) no 

- any problems with kitchen equipment, including cleaning supplies? Oven burnt middle of cake. 

- any cooking questions you need help with? no 

- any other thoughts or suggestions? The burner in the oven is right down the center. The middle of the cake was burnt. The outside of the cake was just fine.